Two solutions of salt water contain 0.04% and 0.19% salt respectively. A lab technician wants to make 1 liter of solution which contains 0.13% salt. How much of each solution should she use?
Answer by mananth(16946)

percent ---------------- quantity
Salt solution I 0.04 ---------------- x L
Salt solution II 0.19 ------ 1 - x L
Mixture 0.13% ---------------- 1

0.04x+0.19(1-x) =0.13* 1

0.04x+0.19-0.19 x=0.13
0.04x- 0.19x =0.13- 0.19
-0.15 x=-0.06
/ -0.15
x= 0.4 L 0.04% Salt solution I
0.6 L 0.19% Salt solution II

m.ananth@hotmail.ca